OpenStudy (anonymous):

Quiz of the day :

OpenStudy (anonymous):

A thin piece of wire 40 meters long is cut into two pieces. One piece is used to form a circle with radius r, and the other is used to form a square. No wire is left over. Which of the following represents the total area, in square meters, of the circular and the square regions in terms of r? (p=> Pi) A. pr2 B. pr2 + 10 C. pr2 + ¼ p2r2 D. pr2 + (40-2pr)2 E. pr2 + (10- ½ pr)2

OpenStudy (jackellyn):

Radius of the circle is r. Circumference of the circle = 2*pi*r Total wire is 40 meters long. Wire used for circle with radius r is 2*pi*r, rest of the wire that is used to form a square will be 40 - 2*pi*r which is nothing but perimeter of square. One side of the square is (40 - 2*pi*r)/4 = 10 - (pi*r)/2 Now area of circle + area of square = pi*r^2 + (10 - (pi*r)/2)^2

OpenStudy (anonymous):

jackellyn good. your answer is correct :)